Opened 9 years ago

Closed 3 years ago

#360 closed flight processing (fixed)

EU10/03, flight day 234a/2010, Lake Balaton (East) Area A

Reported by: anch Owned by: knpa
Priority: alpha 5 Milestone:
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tipo)

Data location: ~arsf/workspace/EU10_03-2010_234a_Lake_Balaton_East_Area_A/

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Lake Balaton provides an ideal area for testing emerging algorithms on retrieval of in-water constituents and on mapping bloom-forming cyanobacteria using in-situ radiometric, airborne and satellite data.

Priority: a5

PI: A. Tyler

Sensors:

  • Eagle (20/10/2010)
  • Hawk (28/02/2011)
  • Leica LIDAR (11/12/2013)
  • Digital Photography (20/10/2010)

Attachments (1)

2342010.sync (111 bytes) - added by iopa 9 years ago.
Eagle SCT offsets

Download all attachments as: .zip

Change History (32)

comment:1 Changed 9 years ago by anch

Please note this is stored in archive-overflow until space becomes available

comment:2 Changed 9 years ago by iopa

Applanix and IPAS have been done.

comment:3 Changed 9 years ago by iopa

Starting eagle and camera processing.

comment:4 Changed 9 years ago by iopa

Camera ready to be delivery checked.

comment:5 Changed 9 years ago by iopa

RCD images left unprocessed are 1-19, as they were pre-flight test images.

comment:6 Changed 9 years ago by emca

Delivery checked camera data from:
~airborne/workspace/EU10_03-2010_234a_Lake_Balaton_East_Area_A/delivery/20101005/EU10_03

No issues with the data - ready to deliver.

comment:7 Changed 9 years ago by iopa

Eagle data has been processed and delivery folder has been created.

Now awaiting a written section in the Readme regarding ASTER dems.

Changed 9 years ago by iopa

Eagle SCT offsets

comment:8 Changed 9 years ago by iopa

Eagle ready for checking.

comment:9 Changed 9 years ago by emca

Delivery check eagle
~airborne/workspace/EU10_03-2010_234a_Lake_Balaton_East_Area_A/delivery/20101008/EU10_03

DEM needs to be added to delivery, and to the readme contents.
Otherwise ready to deliver.

comment:10 Changed 9 years ago by emca

Eagle and Camera sent to A Tyler 20/10/2010

comment:11 Changed 9 years ago by emca

  • Description modified (diff)

comment:12 Changed 9 years ago by benj

  • Description modified (diff)

comment:13 Changed 9 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Started Hawk processing.

comment:14 Changed 9 years ago by knpa

Found a problem with the line counter which needed to be fixed before hawk would process in az properly. Have written a script to overwrite rubbish line counters with a synthetic one, can therefore now continue with hawk processing.

comment:15 Changed 9 years ago by knpa

Lines 7 and 8 have turns.

comment:16 Changed 9 years ago by knpa

Hawk SCTs

1 0.00
2 0.00
3 -0.02
4 0.03
5 0.00
6 0.00
7 0.00
8 0.00
9 0.00
10 0.01
11 0.00

comment:17 Changed 9 years ago by knpa

CORRECTION: Hawk lines 8 and 9 are the ones with turns in them.

Omitted lines 0-1700 from 8
Omitted lines 0-3555 from 9

comment:18 Changed 9 years ago by knpa

Submitted Hawk for delivery check.

comment:19 Changed 9 years ago by emca

Delivery check complete:

Need to chmod the lev1 bil files
Mention in readme that the first ~10 & last ~15 band of hawk are data are not good quality

Otherwise ready to deliver

comment:20 Changed 9 years ago by knpa

Project on disk (97) awaiting the addition of LIDAR data from day 232b as that is also for Dr Tyler.

comment:21 Changed 9 years ago by emca

  • Description modified (diff)

Hawk data sent 28/02/2011

comment:22 Changed 7 years ago by knpa

Delivery check of HS reprocessing complete,

All looks fine apart form:

  • Incorrect file name format on DEM + hdr

comment:23 Changed 6 years ago by tipo

LiDAR delivery created. There is large amounts of noise, large areas of very low intensity, and height offsets between lines sometimes reaching 3 or 4 meters.

comment:24 Changed 6 years ago by stgo

Beginning LiDAR D/C

comment:25 Changed 6 years ago by stgo

D/C complete, only changes to Readme required:

  • data quality needs full stop after "very low intensity"
  • it would be worth noting that lines 01 02 and 03 (017 016 and 015) are considerably worse for overlap than the others.
  • needs updating to have correct UTM grid number.

Otherwise ready to deliver.

comment:26 Changed 6 years ago by tipo

Readme issues fixed.
Made avalable on ftp site (11/12/2013)

comment:27 Changed 6 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

comment:28 Changed 6 years ago by knpa

Merged and deleted workspace version

comment:29 Changed 3 years ago by dac

Started uploading camera and raw data to NEODC. Hyperspectral and LiDAR already available from http://browse.ceda.ac.uk/browse/neodc/arsf/2010/EU10_03/EU10_03-2010_234a_Lake_Balaton_East_Area_A

comment:30 Changed 3 years ago by dac

Data uploaded to NEODC ready for archiving.

comment:31 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Camera and raw data available from NEODC. Marking as closed.

Note: See TracTickets for help on using tickets.